# Random Exclusion Codes: Quantum Advantages of Single-Shot Communication

### v2

@@ Appendix A Optimal success with classical resource in $(2,3)$ REC
Appendix B Maximum quantum success probability in $(2,3)$ REC
Proposition 2.
The maximum average success probability in the quantum random exclusion task for a two-outcome POVM is $\frac{7+\sqrt{2}}{9}$.
Proof.
Without loss of generality, we consider the two 2-outcome POVMs applied in the protocol are: $M_{1}\equiv M_{10}+M_{11}=\mathbb{I}$  (we take $M_{12}=0$) and $M_{2}\equiv M_{20}+M_{21}=\mathbb{I}$  (we take $M_{22}=0$). Let us partition the set of all words $a=a_{1}a_{2}\in\mathcal{A}^{2}$ into two parts: =$\mathbb{P}=\{00,01,10,11\}$ and $\mathbb{P}^{c}=\{02,20,12,21,22\}$. Then, from the Eq. (7) we get
${p}^{\mathrm{(REC)}}_{Q}$ $\!=\!\frac{1}{18}\sum_{a}\sum_{i=1}^{2}{\rm Tr}[(\mathbb{I}-M_{ia_{i}})\rho_{a_{1}a_{2}}]$
$=\frac{4}{9}\left\{\frac{1}{8}\sum\limits_{a\in\mathbb{P}}\sum\limits_{i=1}^{2}{\rm Tr}\left[(\mathbb{I}-M_{ia_{i}})\rho_{a_{1}a_{2}}\right]\right\}$
$+\frac{1}{18}\sum\limits_{a\in\mathbb{P}^{c}}\left\{\sum\limits_{i=1}^{2}{\rm Tr}[(\mathbb{I}-M_{ia_{i}})\rho_{a_{1}a_{2}}]\right\}$
$\leq\frac{4}{9}\left\{\frac{1}{8}\sum\limits_{a\in\mathbb{P}}\sum\limits_{i=1}^{2}{\rm Tr}\left[(\mathbb{I}-M_{ia_{i}})\rho_{a_{1}a_{2}}\right]\right\}+\frac{5}{9}$
$\leq\frac{4}{9}\left\{\max\limits_{\{\begin{subarray}{c}M_{ia_{i}},\\ \rho_{a_{1}a_{2}}\end{subarray}\}}\leavevmode\nobreak\ \frac{1}{8}\sum\limits_{a\in\mathbb{P}}\sum\limits_{i=1}^{2}{\rm Tr}\left[(\mathbb{I}-M_{ia_{i}})\rho_{a_{1}a_{2}}\right]\right\}\!+\!\frac{5}{9}$
$=\frac{4}{9}\left\{\max\limits_{\{\begin{subarray}{c}\tilde{M}_{ia_{i}},\\ \rho_{a_{1}a_{2}}\end{subarray}\}}\leavevmode\nobreak\ \frac{1}{8}\sum\limits_{a\in\mathbb{P}}\sum\limits_{i=1}^{2}{\rm Tr}\left[\tilde{M}_{ia_{i}}\leavevmode\nobreak\ \rho_{a_{1}a_{2}}\right]\right\}+\frac{5}{9}$
(we $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (\mbox{we\leavevmode\nobreak\ rewrite}\leavevmode\nobreak\ \mathbb{I}-M_{ia_{i}}=\tilde{M}_{ia_{i}}\leavevmode\nobreak\ ),$
$\leq\frac{4}{9}\left\{\frac{1}{2}\left(1+\frac{1}{\sqrt{2}}\right)\right\}+\frac{5}{9}=\frac{7+\sqrt{2}}{9}.$ (16)
In the last step above, we simply apply the known result, i.e., in a $(2,2)$ RAC, the quantum maximum for the average success probability is $(1+1/\sqrt{2})/2$. Since our quantum protocol reaches the derived quantum upper bound, we find an optimal over the set of all two-outcome measurement protocols.
In general, the optimization problem consists of maximizing the figure of merit over all possible three-outcome POVMs on a qubit space. To find optimal states and measurements for the purpose, let us rewrite the average success probability as follows,
${p}^{\mathrm{(REC)}}_{Q}$ $=$ $\frac{1}{36}\sum_{a}{\rm Tr}[\sum_{i=1}^{2}[(\mathbb{I}-M_{ia_{i}})\otimes\rho_{a_{1}a_{2}}^{T})|\phi^{+}\rangle\langle\phi^{+}|],$
where $|\phi^{+}\rangle=(|00\rangle+|11\rangle)/\sqrt{2}$. Note that a state $|\phi^{+}\rangle$ is permutationally invariant, which means that the states and the measurements that maximise the success probability should be in a symmetric subspace, supported by a quantum $2$-design. This implies that optimal states and measurements fulfil the condition,
$\sum_{i=1}^{2}(\mathbb{I}-M_{ia_{i}})\propto\rho_{a_{1}a_{2}}^{*}.$ (17)
Applying the above, we will then show that to achieve the maximal success probability, it is sufficient to optimize a function which depends only on the measurement parameters (i.e., for any given decoding measurements, one can always choose an optimal set of states derived from the measurement). All these simplify the construction of states and measurements.
First, recall that we need to maximize ${p}^{\mathrm{(REC)}}_{Q}$ in Eq. (7) in the main text. From the objective function, one can see that the maximum is achieved with extremal (pure) states and measurements. Then let us consider a set of nine pure qubit states
$\rho_{a_{1}a_{2}}$ $=$ $\frac{1}{2}(\mathbb{I}+\hat{n}_{a_{1}a_{2}}\cdot\vec{\sigma}),$ (18)
where $\hat{n}_{a_{1}a_{2}}$ for $a_{1},a_{2}\in\{0,1,2\}$ are unit vectors on the Bloch sphere, and two measurements corresponding to extremal POVMs which are given by rank one projectors, i.e.,
$M_{1a_{1}}$ $=$ $r_{1a_{1}}\left(\frac{\mathbb{I}+\hat{m}_{1a_{1}}\cdot\vec{\sigma}}{2}\right),$ (19)
$M_{2a_{2}}$ $=$ $r_{2a_{2}}\left(\frac{\mathbb{I}+\hat{m}_{2a_{2}}\cdot\vec{\sigma}}{2}\right),$ (20)
where $\hat{m}_{1a_{1}}$ ($\hat{m}_{2a_{2}}$) are unit vectors on the Bloch sphere in a plane passing through the origin such that weights $0\leq r_{1a_{1}}\leq 1$ ($0\leq r_{2a_{2}}\leq 1$) and $\sum_{a_{1}}r_{1a_{1}}=2$  ($\sum_{a_{2}}r_{2a_{2}}=2$). On rearranging the terms in ${p}^{\mathrm{(REC)}}_{Q}$ in Eq. (7), the expression for quantum success probability becomes
${p}^{\mathrm{(REC)}}_{Q}$ $=$ $1-\frac{1}{18}\sum_{a}{\rm Tr}\left[(M_{1a_{1}}+M_{2a_{2}})\rho_{a_{1}a_{2}}\right]$
$=$ $\frac{2}{3}-\frac{1}{36}\sum_{a}\left\{\hat{n}_{a_{1}a_{2}}\cdot(r_{1a_{1}}\hat{m}_{1a_{1}}+r_{2a_{2}}\hat{m}_{2a_{2}})\right\},$
then application of Eq. (17) gives
${p}^{\mathrm{(REC)}}_{Q}$$\leq$ $\frac{2}{3}+\frac{1}{36}\leavevmode\nobreak\ \sum_{a}|r_{1a_{1}}\!\hat{m}_{1a_{1}}\!+\!r_{2a_{2}}\hat{m}_{2a_{2}}|.$ (21)
Note that very similar to the the considered REC task, on analysing the corresponding RAC, through Eq. (2), for maximum quantum success, we get an achievable upper bound function as follows:
${p}^{\mathrm{(RAC)}}_{Q}$$\leq$ $\frac{1}{3}+\frac{1}{36}\leavevmode\nobreak\ \sum_{a}|r_{1a_{1}}\!\hat{m}_{1a_{1}}\!+\!r_{2a_{2}}\hat{m}_{2a_{2}}|.$ (22)
Thus, the optimization problems both for the quantum REC and RAC reduces to maximizing the same objective function.
From the steps above one can easily see that for arbitrarily fixed measurements, one can always choose states such that the upper bound is attained. Therefore, it is sufficient to maximize the quantity
$f\!$ $=$ $\!\sum_{a_{2}\!=\!0}^{2}\!\sum_{a_{1}\!=\!0}^{2}|r_{1a_{1}}\hat{m}_{1a_{1}}\!+\!r_{2a_{2}}\hat{m}_{2a_{2}}|,$
$=$ $\!\sum_{a_{2}\!=\!0}^{2}\!\sum_{a_{1}\!=\!0}^{2}\!\!\left\{r_{1a_{1}}^{2}\!\!+\!r_{2a_{2}}^{2}\!\!+\!2r_{1a_{1}}\!r_{2a_{2}}\!\cos(\phi_{1a_{1}}\!\!-\!\phi_{2a_{2}})\sin\theta_{2a_{2}}\right\}^{\frac{1}{2}},$
which is a function of the two measurements derived from the extremal POVMs $M_{i}$, where $i\in\{1,2\}$ and $(\theta_{ia_{i}},\phi_{ia_{i}})$ are spherical polar coordinates of the Bloch vector $\hat{m}_{ia_{i}}$. The weight parameters of the three-outcome (corresponding to indices $\kappa\in\{0,1,2\}$) POVMs, can be written by
$r_{i\kappa}\!$ $=$ $\!\frac{2\sin\alpha_{i\kappa}}{\sin\alpha_{i0}\!+\!\sin\alpha_{i1}\!+\!\sin\alpha_{i2}},$ (24)
$\alpha_{i0}+\alpha_{i1}+\alpha_{i2}=2\pi,$ (25)
$0\leq\alpha_{i0},\leavevmode\nobreak\ \alpha_{i1},\leavevmode\nobreak\ \alpha_{i2}\leq\pi.$ (26)
See Fig. 3 to visualize the measurements.
Figure 3: Illustration of a three-outcome extremal POVM (on a qubit) on a plane passing through the origin of the Bloch sphere.
In general, the Bloch vectors corresponding to two different measurements may live in two different planes. We choose, without loss of generality, the Bloch plane of the measurement $M_{1}$ as the $XY$-plane. We choose the $Y$-axis along the line where the planes of the two measurements $M_{1}$ and $M_{2}$ intersect. Let us denote angle of the unit vector normal to the plane of the measurement $M_{2}$ with the $Z$-axis by $0\leq\Theta\leq\pi$. Further, without loss of generality, for the second measurement $M_{2}$ we set the angle of the Bloch vector of the first outcome $\phi_{20}=0$, and for the first measurement let us denote the angle of the Bloch vector of the first outcome $\phi_{10}=\Phi$. Now the objective function in Eq. (LABEL:eq:f) can be re-written in terms of the variables $\alpha_{ia_{i}}$, $\Theta$, and $\Phi$ by applying Eq. (24) and transforming the variables $(\theta_{ia_{i}},\phi_{ia_{i}})$ to the new variables.
[…]
$\hat{m}_{2j}=(\sin\theta_{2j}\cos\phi_{2j},\leavevmode\nobreak\ \sin\theta_{2j}\sin\phi_{2j},\leavevmode\nobreak\ \cos\theta_{2j})$ (27)
One can solve the system of equations
$\hat{\mathbf{n}}_{2}\cdot\hat{m}_{2j}$ $=$ 00 0 (28)
$\hat{\mathbf{p}}_{2}\cdot\hat{m}_{2j}$ $=$ $\cos\alpha_{2j}$ (29)
to find the variables $\theta_{2j}$ and $\phi_{2j}$ in terms of the variable $\Theta$ and $\alpha_{2j}$. Thus, the objective function $f$ in Eq. (LABEL:eq:f) is now written in terms of six free variables: i.e., $f(\Theta,\Phi,\alpha_{10},\alpha_{12},\alpha_{20},\alpha_{22})$ (here we have applied Eq. (25) to eliminate the two dependent variables $\alpha_{11}$ and $\alpha_{21}$). Then the following proposition holds.
Proposition 3.
The optimal value of $f$ in Eq. (LABEL:eq:f) is achieved with both the three outcome POVMs $M_{1}$ and $M_{2}$ in the $XY$-plane of the Bloch sphere, that is, when variable $\Theta=0$.
Proof.
—We omit the lengthy expressions and steps, but the idea is simple. Objective function in Eq. (LABEL:eq:f) is written in terms of $\Theta$ and the remaining variables. Then we analyze the effect of changing the angle $\Theta$ (angle between the two measurement planes). This is done by taking the partial derivative of the objective function $f$ with respect to the variable $\Theta$. We find that for an arbitrarily fixed value of all the other variables, the partial derivative with respect to $\Theta$ is either always nonnegative or always nonpositive, i.e., the function is either increasing or decreasing. Then a maximum (for any fixed value of all the other variables) will be achieved at an end point in the interval $0\leq\Theta\leq\pi$, implying that the optimal measurement $M_{2}$ lies in the $XY$-plane of the Bloch sphere (i.e., both the measurements are in the same plane). ∎
Applying the above lemma, now the optimization problem is reduced to maximizing
$\!\!\!\!\!\!f^{\ast}\!$ $=$ $\!\!\sum_{a_{2}\!=\!0}^{2}\!\sum_{a_{1}\!=\!0}^{2}\!\!\left\{\!r_{1a_{1}}^{2}\!\!+\!r_{2a_{2}}^{2}\!\!+\!2r_{1a_{1}}\!r_{2a_{2}}\!\cos(\phi_{1a_{1}}\!\!-\!\phi_{2a_{2}})\right\}^{\frac{1}{2}\!}$
where both the three outcome measurements are from the $XY$-plane. In Fig. (4) two three outcome extremal POVMs in the XY-plane is shown.
Figure 4: Illustration of three outcome extremal POVMs on the $XY$-plane of the Bloch sphere.
In Eq. (LABEL:eq:fstar), for $i\in\{1,2\}$, we substitute
$r_{i0}$ $=$ $\frac{2\sin\alpha_{i0}}{\sin\alpha_{i0}-\sin(\alpha_{i0}+\alpha_{i2})+\sin\alpha_{i2}},$ (31)
$r_{i1}$ $=$ $\frac{-2\sin(\alpha_{i0}+\alpha_{i2})}{\sin\alpha_{i0}-\sin(\alpha_{i0}+\alpha_{i2})+\sin\alpha_{i2}},$ (32)
$r_{i2}$ $=$ $\frac{2\sin\alpha_{i2}}{\sin\alpha_{i0}-\sin(\alpha_{i0}+\alpha_{i2})+\sin\alpha_{i2}};$ (33)
$\phi_{i0}$ $=$ $(2-i)\leavevmode\nobreak\ \Phi,$ (34)
$\phi_{i1}$ $=$ $(2-i)\leavevmode\nobreak\ \Phi+\alpha_{i2},$ (35)
$\phi_{i2}$ $=$ $(2-i)\leavevmode\nobreak\ \Phi+\alpha_{i2}+\alpha_{i0}.$ (36)
The free variables in the optimization problem are now explicit; constraints on these variables are given by
$0\leq\alpha_{i0},\leavevmode\nobreak\ \alpha_{i2}\leq\pi\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \pi\leq\alpha_{i0}+\alpha_{i2}\leq 2\pi,$ (37)
$0\leq\Phi\leq\pi.$ (38)
Proposition 4.
The optimal value of $f^{\ast}$ in Eq. (LABEL:eq:fstar) is given by $4(1+\sqrt{2})$ and it is achieved when both measurements $M_{1}$ and $M_{2}$ are two outcome projective measurements in orthogonal directions. Consequently, the optimal quantum success probability for the REC task is $(7+\sqrt{2})/9$ (and that for the corresponding RAC task is $(4+\sqrt{2})/9$); in particular, the quantum protocol presented in the main text is an optimal one for the REC task.
Our proposition is supported by our analytical simplification (reduction) of the optimization problem stated in proposition (3), followed by numerical optimization implemented on Mathematica. We get an approximate maximum value $\max(f^{\ast})\approx 9.65685\approx 4(1+\sqrt{2})$.
Appendix C Bounds on random exclusion task for alphabet size m
Proposition 5.
The maximum classical average success probability for the random exclusion task with alphabet size $m$ is $\max_{C}\leavevmode\nobreak\ {p}^{\mathrm{(REC)}}=(1-\frac{1}{m^{2}})$.
Proof.
Let us first find the optimal value over all the deterministic protocols. The two stages of any deterministic protocol are as follows. (i) Alice sending one bit information $c\in\{0,1\}$ corresponds to some two partition $\mathbb{P}_{0},\mathbb{P}_{1}$ of the set of all possible words $\mathcal{A}^{2}:=\{a_{1}a_{2}:a_{1},a_{2}\in\{0,1,2,\dots,m-1\}\}$; if the received word $a=a_{1}a_{2}\in\mathbb{P}_{0}$, Alice communicates $c=0$ to Bob, otherwise she sends $c=1$. (ii) Bob’s answer depends on the received bit, he answers with, say, $b_{i}^{c}\in\{0,1,2,\dots,m-1\}$, i.e., Bob’s deterministic strategy corresponds to assigning values to $b_{1}^{0},b_{2}^{0},\leavevmode\nobreak\ b_{1}^{1},b_{2}^{1}\in\{0,1,2,\dots,m-1\}$.
First we show that for any deterministic strategy of Alice and Bob, for at least two out of the total $2m^{2}$ possible instances ($m^{2}$ possible words to Alice and two possible questions to Bob), Bob’s answers will be incorrect. To show this, first, we note that from any deterministic strategy of Bob one can derive the following two words $b_{1}^{0}b_{2}^{0},b_{1}^{1}b_{2}^{1}\in\mathcal{A}^{2}$, therefore, for these two words, given a deterministic strategy of Alice, there are four possibilities:
  1. 1.
    $b_{1}^{0}b_{2}^{0},\leavevmode\nobreak\ b_{1}^{1}b_{2}^{1}\in\mathbb{P}_{0}$,
  2. 2.
    $b_{1}^{0}b_{2}^{0},\leavevmode\nobreak\ b_{1}^{1}b_{2}^{1}\in\mathbb{P}_{1}$,
  3. 3.
    $b_{1}^{0}b_{2}^{0}\in\mathbb{P}_{0}$ and $b_{1}^{1}b_{2}^{1}\in\mathbb{P}_{1}$,
  4. 4.
    $b_{1}^{0}b_{2}^{0}\in\mathbb{P}_{1}$ and $b_{1}^{1}b_{2}^{1}\in\mathbb{P}_{0}$.
In the case-1, when Alice receives the word $b_{1}^{0}b_{2}^{0}$, Bob’s answer is $b_{i}^{0}$ which is incorrect at both the positions $i\in\{1,2\}$. In case-2, similar to case-1, when Alice receives the word $b_{1}^{1}b_{2}^{1}$, Bob’s answer is $b_{i}^{1}$ which is incorrect at both the positions $i\in\{1,2\}$.
In case-3 and case-4 it can never happen that $b_{1}^{0}b_{2}^{0}=b_{1}^{1}b_{2}^{1}$ because same word cannot appear in two distinct partitions.
For the case-3, when Alice receives the word $b_{1}^{0}b_{2}^{0}$ ($b_{1}^{1}b_{2}^{1}$), Bob’s answer is $b_{i}^{0}$ ($b_{i}^{1}$) which is incorrect at both the positions $i\in\{1,2\}$; thus in the case-3 Bob answers incorrectly in at least four instances, namely $b_{1}^{0}b_{2}^{0},b_{1}^{1}b_{2}^{1}$.
Lastly, in the case-4, first suppose that $b_{1}^{0}b_{2}^{1}\neq b_{2}^{0}b_{1}^{1}$. Then for the word $b_{1}^{0}b_{2}^{1}$, to whichever partition it belongs, Bob’s answer will be incorrect at least at one position $i\in\{1,2\}$; similarly, for the word $b_{2}^{0}b_{1}^{1}$, to whichever partition it belongs, Bob’s answer will be incorrect at least at one position $i\in\{1,2\}$. In the remaining situation, i.e., $b_{1}^{0}b_{2}^{1}=b_{2}^{0}b_{1}^{1}\Leftrightarrow b_{1}^{0}=b_{2}^{0}\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ b_{1}^{1}=b_{2}^{1}$, on considering the words $b_{1}^{0}b_{1}^{1}$ and $b_{1}^{1}b_{1}^{0}$ (irrespective which partition they belong to) Bob’s answer is incorrect at least two times; at least once for the first word $b_{1}^{0}b_{1}^{1}$ and at least once for the second word $b_{1}^{1}b_{1}^{0}$. Then, in case-4 as well, Bob’s answer is incorrect in least in two instances.
Therefore, in all the four cases, we get that for any deterministic strategy of Alice and Bob, in at least two out of $2m^{2}$ cases Bob answers incorrectly. Therefore, maximal average success probability with deterministic strategies is upper bounded by $\frac{2m^{2}-2}{2m^{2}}=1-\frac{1}{m^{2}}$.
Finally, we show that probabilistic strategies cannot do better. Lets say there are total $N$ number of deterministic strategies $S_{j}$ where $j\in\{1,\cdots,N\}$. Suppose, Alice and Bob have shared random variables $\{\lambda_{j}\}_{j=1}^{N}$ with probability distribution $p(\lambda_{j})=p_{j}$. Note that any probabilistic strategy can be implemented with an access to the considered shared random variables. Then if Alice and Bob apply strategy $S_{j}$ with probability $p_{j}$ we get a probabilistic strategy $S=\{p_{j},S_{j}\}_{j=1}^{N}$. Then the average success probability with any such strategy $S$ is ${p}^{\mathrm{(REC)}}_{S}=\sum_{j}p_{j}{p}^{\mathrm{(REC)}}_{S_{j}}\leq\sum_{j}p_{j}\left(1-\frac{1}{m^{2}}\right)=1-\frac{1}{m^{2}}$. This completes the proof. ∎
Quantum advantage.— Here we will show that there is a quantum protocol which considers only two outcome measurements for decoding to beat the optimal classical value derived above. Alice sends the information on her words $a_{1}a_{2}$ by encoding in a one qubit state $\rho_{a_{1}a_{2}}$ as follows:
$\rho_{00}$ $=\frac{1}{2}\left(\mathbb{I}+\sigma_{x}\right),$ $\rho_{01}$ $=\frac{1}{2}\left(\mathbb{I}+\sigma_{z}\right),$ $\rho_{0c_{2}}$ $=\frac{1}{2}\left(\mathbb{I}+\frac{\sigma_{x}+\sigma_{z}}{\sqrt{2}}\right)$
$\rho_{10}$ $=\frac{1}{2}\left(\mathbb{I}-\sigma_{z}\right),$ $\rho_{11}$ $=\frac{1}{2}\left(\mathbb{I}-\sigma_{x}\right),$ $\rho_{1c_{2}}$ $=\frac{1}{2}\left(\mathbb{I}-\frac{\sigma_{x}+\sigma_{z}}{\sqrt{2}}\right)$
$\rho_{c_{1}0}$ $=\frac{1}{2}\left(\mathbb{I}+\frac{\sigma_{x}-\sigma_{z}}{\sqrt{2}}\right),$ $\rho_{c_{1}1}$ $=\frac{1}{2}\left(\mathbb{I}+\frac{\sigma_{z}-\sigma_{x}}{\sqrt{2}}\right),$ $\rho_{c_{1}c_{2}}$ $=\frac{\mathbb{I}}{2}$
where $\text{where }c_{1},c_{2}$$\in\{2,3,\ldots,m-1\}.$ (39)
All the qubit states used for encoding are from the $XZ$-plane of the Bloch sphere. To answer the first position in the Alice’s word, Bob performs a two outcome projective measurement $\mathcal{M}_{1}$; on measurement outcome $+1$ ($-1$) he answers 00 ($1$). Whereas for answering the second position he measures $\mathcal{M}_{2}$ and answers with 00 ($1$) when the measurement outcome is $+1$ ($-1$). The average success probability of our protocol is
${p}^{\mathrm{(REC)}}_{Q}$ ::=:= $1-\frac{1}{2m^{2}}\sum_{a_{1},a_{2}\in\mathcal{A}}\leavevmode\nobreak\ \mbox{Tr}\left[\rho_{a_{1}a_{2}}\left(M_{1a_{1}}+M_{2a_{2}}\right)\right]$
$=$ $1-\frac{2-\sqrt{2}}{m^{2}}.$
Here the first and second measurements are implemented as follows
$\mathcal{M}_{1}$$\equiv$ $\left\{M_{10}=\frac{1}{2}\left(\mathbb{I}-\frac{\sigma_{x}+\sigma_{z}}{\sqrt{2}}\right),\leavevmode\nobreak\ M_{11}=\frac{1}{2}\left(\mathbb{I}+\frac{\sigma_{x}+\sigma_{z}}{\sqrt{2}}\right),\leavevmode\nobreak\ M_{1c_{1}}=0\right\},$
$\mathcal{M}_{2}$$\equiv$ $\left\{M_{20}=\frac{1}{2}\left(\mathbb{I}+\frac{\sigma_{z}-\sigma_{x}}{\sqrt{2}}\right),\leavevmode\nobreak\ M_{21}=\frac{1}{2}\left(\mathbb{I}+\frac{\sigma_{x}-\sigma_{z}}{\sqrt{2}}\right),\leavevmode\nobreak\ M_{2c_{2}}=0\right\}.$ (40)
Note that Bob never answers $\{2,3,\ldots,m-1\}$, still our protocol gives a high success because the goal here is to answer with alphabet different from those appearing in the words. Also note that the state $\rho_{c_{1}c_{2}}$ used for encoding the words $c_{1}\;c_{2}$, can be any arbitrary state, even any mixed state, e.g., maximally mixed state $\frac{\mathbb{I}}{2}$. This is due to the fact $M_{1c_{1}}=0$ and $M_{2c_{2}}=0$, i.e., outcomes $c_{1}\text{ or }c_{2}$ never occur for the considered measurements.
Proposition 6.
The maximum average success probability in the quantum random exclusion task, with alphabet size $m$, for a two-outcome POVM is achieved with the above discussed quantum protocol.
Proof.
Let us address the scenario wherein Alice sends a qubit instead of a bit. We can show that there is a quantum advantage in this case for all $m$. The proof of advantage only requires Bob to perform two 2-outcome measurements. Say we form the partition as =$\mathbb{Q}=\{00,01,10,11\}$ and ¯=$\overline{\mathbb{Q}}=\{a_{1}a_{2}\in\mathcal{A}^{2}:a_{1}a_{2}\notin\mathbb{Q}\}$. The two 2-outcome measurements can be used to distinguish between states associated with $\mathbb{Q}$. For the states associated with $\overline{\mathbb{Q}}$, the measurement output is always correct for our task. We can find an upperbound for the quantum success probability as well:
${p}^{\mathrm{(REC)}}_{Q}$ $=\frac{1}{2m^{2}}\sum_{a\in\mathcal{A}^{2}}\sum_{i=1}^{2}{\rm Tr}\left[(\mathbb{I}-M_{ia_{i}})\rho_{a_{1}a_{2}}\right]$
[…]

### v3

@@ IV.1 Optimal bit implementation
IV.2 Qubit implementation and quantum advantage
We next present a $(2,3)$ REC with single qubit states in the following,
where
$|\psi_{02}\rangle$ $=$ $\frac{|0\rangle+|1\rangle}{\sqrt{2}},\leavevmode\nobreak\ |\psi_{12}\rangle=\frac{|0\rangle-|1\rangle}{\sqrt{2}},\leavevmode\nobreak\ \mathrm{and}\leavevmode\nobreak\ \rho_{22}=\mathbb{I}/2.$
We construct a decoding with two-outcome measurements,
$M_{1}=\{M_{10},M_{11},M_{12}\}\leavevmode\nobreak\ \mathrm{and}\leavevmode\nobreak\ M_{2}=\{M_{20},M_{21},M_{22}\}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\$(8)
where
and $|\pm\rangle=(|0\rangle\pm|1\rangle)/\sqrt{2}$ and $|\pm i\rangle=(|0\rangle\pm i|1\rangle)/\sqrt{2}$. Notice that the states and measurements above are on a great circle of the Bloch sphere. We display the states and measurements of our scheme in in Fig. 3.
Figure 3: Illustration of states and measurements for the scheme discussed in text. States are displayed in light blue and the measurements $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ in red and green respectively.
The quantum protocol works as follows. Alice sends a state $\rho_{a_{1}a_{2}}$ and Bob first decides which one, $a_{1}$ or $a_{2}$, he is going to read. For the first trit $a_{1}$, he applies measurement $M_{1}$, and for the second, $a_{2}$, he applies $M_{2}$. The probability of successfully excluding a one-bit message of Alice is given by,
$p^{\mathrm{(REC)}}_{Q}$ $=$ $\frac{1}{18}\sum_{a}\sum_{i=1}^{2}{\rm Tr}[(M_{i(a_{i}\oplus 1)}+M_{i(a_{i}\oplus 2)})\leavevmode\nobreak\ \rho_{a_{1}a_{2}}]\leavevmode\nobreak\ \leavevmode\nobreak\$(9)
where $\oplus$ represents addition modulo three. It follows that
$p^{\mathrm{(REC)}}_{Q}=\frac{7+\sqrt{2}}{9}$ (10)
which is larger than the maximum classical value $8/9$ (Proposition 1.). Thus, the utility of quantum resources in RECs is demonstrated.
IV.3 Maximum quantum advantage
There may exist states and measurements other than those in Eqs. (7) and (8), in particular using POVMs containing no null elements, such that the success probability is even higher. That is, one can ask if the bound in Eq. (10) is maximal. Our answer is in the affirmative. Firstly, we find that to maximize the success probability with qubit transmission, it suffices to restrict states and measurements to a half-plane on a Bloch sphere. This also holds true even if parties aim at a guessing task, and hence, applies to RACs with qubits. Secondly, extensive numerical optimization of the average success probability finds Eq. (10) with two POVM elements are zeros. We detail the calculations in Appendix LABEL:appendA. The gap between quantum and classical bounds is
which is approximately $0.046$.
For comparison, let us also consider a $(2,3)$ RAC with single-qubit transmission. An upper bound to success probabilities with classical strategies is known as $5/9$ deba+PRA2023. We compute the maximum average probability with quantum resources, applying the aforementioned result of reducing to a great circle. We find that a relabelling of our quantum protocol for the REC gives the quantum maximum for the RAC (see Appendix LABEL:appendA); messages $a=a_{1}a_{2}$ are encoded in the states orthogonal to ones given in Eq. (7) and the same measurements as in Eq. (8) are applied for decoding. The quantum bound is obtained as $(4+\sqrt{2})/9$. Interestingly, the gap is identical to Eq. (11), i.e.,
$\Delta_{QC}^{(\mathrm{RAC})}=\Delta_{QC}^{(\mathrm{REC})}$ (12)
for the considered $(2,3)$ RAC and $(2,3)$ REC.
To summarise, we have characterised bit and qubit implementations of the $(2,3)$ REC and shown that using quantum systems allows for a higher average probability of success. We also contrasted the behaviour of the $(2,3)$ RAC.
V Dimensional Advantages
In what follows, we are motivated by the observation that a single-shot scenario gives one outcome among possible detection events. This raises the question of which dimensional resources are required to observe distinctions between classical and quantum behaviours in the single-shot setting. Having characterised above the optimal performance of the $(2,3)$ REC in two-dimensional systems, we now turn our attention to a more general setting. We investigate the minimal quantum and classical resources to describe possible detection events in an errorless $(2,3)$ REC. We first note that there is no difference in behaviour for the standard scenario. This further motivates us to add an extra constraint, uniformity of outcomes, to our protocol. Classical implementations of this modified protocol require nine-dimensional systems, whereas quantum systems of four-dimensional behaviour, showing a dimensional advantage in uniform RECs. This behaviour is contrasted with RACs, where no such advantage can be observed.
V.1 Perfect random exclusion codes
We first note that no dimensional advantage can be observed for zero-error implementations of the basic $(2,3)$ REC which has been discussed so far. To do this, we note that there exists a three-dimensional classical, and therefore also quantum, strategy which may implement this task with no error. The explicit strategy is for Alice to encode her messages using the partition
$\begin{split}\mathbb{P}_{0}&=\{11,12,22\}\\ \mathbb{P}_{1}&=\{00,02,20,22\}\\ \mathbb{P}_{2}&=\{01,10\}\end{split}$ (13)
and send the relevant trit value to Bob. If he announces the value he has received, he will exclude without error. As we saw in the previous section, there exists neither a classical nor a quantum strategy that is able to achieve an average success probability of one using a two dimensional implementation. We can thus conclude that three is the minimal dimension requirement in both cases.
V.2 Uniform random exclusion codes
In order to witness a dimensional advantage in the $(2,3)$ REC, we must therefore analyse a modified form of the problem. This extra requirement will be uniformity of the outcomes, meaning that the correct answers by Bob all occur with equal probability. This constraint is regularly considered in the literature on exclusion tasks (see, e.g., Ref Heinosaari2024Simple; PhysRevResearch.2.033374) in order to avoid certain trivial solutions to the problem of state exclusion. It is in fact uniform state exclusion that can be considered a strong signature of quantum, rather than classical, behaviour. The constraint can be furthermore justified if we consider that Alice and Bob are trying to convince a referee, who provides Alice with her string, of their ability to perform quantum exclusion. If their results are non-uniform, then the referee will be able to use their strategy against them when the task is performed over several rounds. For example, in our optimal qubit scheme, Bob will never announce two as his answer. The referee can thus always send strings that do not contain this value, lowering their probability of success. To summarise, our aim here will be to determine the minimum quantum and classical dimensions required to implement a zero-error, uniform $(2,3)$ REC.
The natural framework for studying dimensional requirements is that of communication matrices Kerppo2023Communication, which we review here for the reader’s convenience. A communication matrix is a row-stochastic matrix encoding the conditional probabilities of prepare-and-measure experiments. If the inputs are a set of classical labels $\alpha\in[0,m-1]$ and the outputs a set of classical labels $\beta\in[0,n-1]$ then the communication matrix is an $m\times n$-sized row-stochastic matrix $C$ in which
$C_{\alpha\beta}={\rm Prob}_{M|P}(\beta|\alpha),$ (14)
where the subscripts $P$ and $M$ denote preparation and measurement respectively. For instance, the communication matrix of a binary state discrimination task in which two states are perfectly distinguished would be the diagonal matrix $S_{2}={\rm diag}[1,1]$. Similarly, a uniform state exclusion task with three states excluded in an unbiased manner would correspond to the communication matrix
$A_{3}=\frac{1}{2}\begin{bmatrix}0&1&1\\ 1&0&1\\ 1&1&0\end{bmatrix}$. (15)
This can be implemented in a quantum system by the trine-antitrine construction given earlier in Eqs. 1 and 2 .
[…]